# Res-Bench: Benchmarking the Robustness of Multimodal Large Language Models to Dynamic Resolution Input

### v1

@@ Abstract
Multimodal Large Language Models (MLLMs) increasingly support dynamic image resolutions. However, current evaluation paradigms primarily assess semantic performance, overlooking the critical question of resolution robustness - whether performance remains stable across varying input resolutions. To address this gap, we introduce Res-Bench, a comprehensive benchmark comprising 14,400 samples across 12 resolution levels and six core capability dimensions. We designed a novel evaluation framework that goes beyond traditional accuracy metrics to capture performance stability. This framework introduces multiple robustness metrics: Spearman’s correlation for assessing resolution-performance trends, and Absolute/Relative Continuous Error (ACE/RCE) for measuring performance volatility. Using these metrics, we conducted a large-scale evaluation of leading MLLMs. Our analysis encompasses: (1) model-centric and task-centric robustness examination, (2) investigation of preprocessing strategies including padding and super-resolution, and (3) exploration of fine-tuning for stability enhancement. Our code is available in the supplementary material.
@@ 1 Introduction
Visual perception and reasoning constitute fundamental components of human perceptual and cognitive capabilities. Notably, humans demonstrate remarkable robustness in processing and interpreting visual information across varying image resolutions. This is largely attributable to the adaptive capabilities of our visual system and the brain’s cognitive completion mechanisms (Marr 2010; Briscoe 2011). Specifically, we can adaptively focus on important regions of an image, and for low-resolution images, our brains can fill in missing details based on past experiences and prior knowledge (Walsh and Kulikowski 1998). The quest to equip machines with such sophisticated perceptual abilities is a central theme in modern AI research. However, while multimodal large language models have demonstrated exceptional capabilities across various domains, it remains unclear how well MLLMs can achieve resolution robustness comparable to human visual systems.
Contemporary advanced MLLMs have incorporated sophisticated designed architectures to enable support for dynamic resolution input. These strategies generally fall into two categories: (1) Native dynamic processing method, such as the Vision Transformer (ViT) (Dosovitskiy et al. 2021) in Qwen2.5-VL (Bai et al. 2025) which employs Multimodal Rotary Position Embedding (MRoPE) and window attention to process images at their native resolution; and (2) Patch-based method, utilized by models like InternVL2.5 (Chen et al. 2024b), LLaVA-OneVision (Li et al. 2024b) and LLaVA-UHD (Guo et al. 2024), which segment high-resolution images into smaller sub-images and thumbnails for individual processing. Current research paradigms predominantly focus on benchmarking model performance at the semantic level, including visual perception (Fu et al. 2024), visual question answering (VQA) (Antol et al. 2015), visual reasoning (Lu et al. 2024), etc.. However, insufficient attention has been directed to the robustness discussion at the visual processing level. For each MLLM, three critical ingredients contribute to the resolution robustness: (1) the processing mechanism for dynamic resolution input entailed in the vision encoder (Liu et al. 2021), (2) the level of question spanning both perceptual and cognitive dimensions (Vaishnav and Tammet 2025), and (3) the distribution of training data (Bai et al. 2024).
@@
We conduct extensive experiments on Res-Bench to evaluate their robustness to visual inputs. Additionally, for the obtained low-resolution images, we employ data preprocessing methods like white-padding and super-resolution, aiming to validate their effectiveness in both enhancing absolute performance and improving resolution robustness. The evaluation covers multiple proprietary models (GPT-4o( (OpenAI 2024), and Gemini-1.5 Pro (Team 2024)) and open-source models (e.g.LLaVA-OneVision (Li et al. 2024b), Qwen2.5-VL (Bai et al. 2025), mPLUG-Owl3 (Ye et al. 2024), InternVL2.5(Chen et al. 2024b), etc.). We find there is a pervasive lack of resolution robustness. To understand the underlying factors contributing to this instability, we conducted an in-depth analysis of the experimental results, revealing several novel findings that pose new challenges for the development of MLLMs:
@@ 2 Related Work
Multimodal Large Language Model. A fundamental challenge in Multimodal Large Language Models (MLLMs) (Bai et al. 2025; Li et al. 2024b; Team 2024) lies in effectively processing visual information, particularly at high resolutions where fine-grained details are critical. Early models addressed this by resizing images to a low, fixed resolution, a method prone to distorting critical details necessary for tasks like OCR or fine-grained perception (Radford et al. 2021; Bai et al. 2023). To address these issues, several common solutions have emerged: (1) Native resolution processing methods, such as those used in Qwen2.5-VL (Bai et al. 2025) and Kimi-VL (Team et al. 2025). (2) Patch-based method: employed by models like LLaVA-OneVision (Li et al. 2024b), InternVL2.5 (Chen et al. 2024b), and DeepSeek-VL2 (Wu et al. 2024). While both approaches handle high-resolution inputs, the trade-offs between them regarding resolution robustness remain largely unexplored. This paper systematically investigates the impact of these distinct architectural choices on model stability.
MLLM Benchmarks. The evaluation of MLLMs has rapidly evolved, moving from single-task benchmarks like VQA (Antol et al. 2015; Bigham et al. 2010), caption (Lin et al. 2014; Plummer et al. 2015) to comprehensive, multi-faceted evaluation suites. Recent efforts have produced a rich ecosystem of benchmarks targeting a wide array of high-level capabilities, including: Visual perception (Fu et al. 2024), Chart reasoning (Masry et al. 2022, 2025; Wang et al. 2024c), Mathematical reasoning (Lu et al. 2024; Zhang et al. 2024), Spatial reasoning (Wang et al. 2024a; Li et al. 2025; Xu et al. 2025), Expert-level multimodal understanding (Yue et al. 2024, 2025), Optical Character Recognition (OCR) capabilities (Liu et al. 2024; Yang et al. 2024). Additionally, comprehensive benchmarks (Li et al. 2023; Chen et al. 2024a) have been developed to evaluate overall multimodal intelligence. While their source images may have varying native resolutions, these benchmarks ignore the impact of the resolution distribution.
Robustness Evaluation. Evaluating the robustness of MLLMs—their ability to maintain stable performance against input perturbations—is a critical area of research. Existing work has primarily focused on several key dimensions: (1) robustness against semantic-level perturbations (Cao et al. 2024; Liu et al. 2025); (2) robustness against stylistic or natural variations in images (Cai et al. 2023; Li et al. 2024a); and (3) robustness against common image corruptions (Croce et al. 2021; Li et al. 2024c). A direct comparison of these approaches, summarized in Table 1, reveals a shared limitation: while they comprehensively examine model stability, they primarily test against content-level corruptions and adversarial inputs. The impact of fundamental visual properties like resolution on model performance remains a critical, under-explored area. We introduce Res-Bench to systematically fill this gap, providing the first dedicated benchmark for evaluating MLLM resolution robustness.
@@ 3.1 Data Collection Process
Preliminary Data Filtering. To establish a comprehensive and high-quality benchmark, we selected 13 existing benchmarks as data sources. These include benchmarks designed for general evaluation of MLLMs (e.g., SEED (Li et al. 2023)) as well as specialized benchmarks for assessing specific capabilities of MLLMs (including AI2D (Kembhavi et al. 2016), HR-Bench (Wang et al. 2024b), NaturalBench (Li et al. 2024a), SpatialEval (Wang et al. 2024a), MMMU (Yue et al. 2024), MMMU-Pro (Yue et al. 2025), ChartQA-Pro (Masry et al. 2025), BLINK (Fu et al. 2024), MathVerse (Zhang et al. 2024), MathVista (Lu et al. 2024), OCRBench (Liu et al. 2024), CCOCR (Yang et al. 2024)). To ensure the data consists exclusively of high-quality images, we first conducted preliminary filtering by selecting all high-resolution images (with both width and height exceeding 1344 px) as the foundational dataset. Building upon this foundation, we further refined the dataset by leveraging powerful models to select samples that meet visual dependency criteria. Specifically, we employed GPT-4o (OpenAI 2024) and InternVL-2.5-72B (Chen et al. 2024b) as evaluators for this filtering process. We conducted text-only input evaluations by providing textual questions without any accompanying images, while employing circular evaluation to mitigate random biases. A sample was automatically discarded if both models could answer it correctly without the image, as this indicated weak visual dependency. After this filtering process, the number of candidate samples was reduced from 69840 to 2868.
Manual Verification. To further ensure data quality, we performed human expert review on the pre-filtered samples. The evaluation criteria remained consistent with the aforementioned standards. Additionally, we further optimized the data distribution to ensure better balance. To comprehensively evaluate the capabilities of large models across diverse question types and facilitate systematic assessment, we further refined the phrasing of questions for a subset of the data. For instance, we restructured the single-image multi-question format in NaturalBench (Li et al. 2024a) to a single-question multi-image paradigm, while also refining certain question formulations in CCOCR (Yang et al. 2024). We then systematically generated the full Res-Bench dataset by creating multiple resolution versions for each image, resulting in a total of 14,400 data instances for evaluation. Our data filtering pipeline is illustrated in Figure 2.
@@ 4 Experiment
Baseline. We use Res-Bench to evaluate several leading MLLMs. This includes two powerful proprietary models, GPT-4o (OpenAI 2024) and Gemini 1.5-Pro (Team 2024), and six powerful open-source models employing different dynamic resolution strategies. Qwen2.5-VL (Bai et al. 2025) , Kimi-VL-Instruct (Team et al. 2025) directly process native high-resolution images, while InternVL-2.5 (Chen et al. 2024b), LLaVA-OneVision (Li et al. 2024b) , MiniCPM-o-2.6 (Hu et al. 2024) use patch-based method. We also tested model that do not support dynamic resolution for single images, mPLUG-Owl3 (Ye et al. 2024).
@@ 4.2 Analysis of Preprocessing Strategies
Restoring Information Content via Super-Resolution. We investigated whether algorithmically restoring visual details could recover lost performance more effectively. We employed DiffIR (Xia et al. 2023), a classical diffusion-based SR model, to upscale low-resolution images (224px and 448px) to higher resolutions (448px and 896px). Unlike padding, SR aims to intelligently “fill in” missing information, enhancing image clarity. As shown in Table 3, SR-enhanced images significantly outperform their original low-resolution counterparts and also surpass the performance of the simple padded images.
@@ A.2 Experiment Details
This section provides a detailed, per-model breakdown of results on the Res-Bench dataset. We evaluated a diverse suite of eight leading MLLMs, including proprietary models like Gemini 1.5 Pro (Team 2024) and GPT-4o (OpenAI 2024), as well as open-source models such as Qwen2.5-VL (Bai et al. 2025), LLaVA-OneVision (Li et al. 2024b), InternVL2.5 (Chen et al. 2024b), DeepSeek-VL(Wu et al. 2024), MiniCPM-o-2.6 (Hu et al. 2024) and mPLUG-Owl3 (Ye et al. 2024). For each of the evaluated models, we present the full results across all 15 sub-tasks, including accuracy at each of the 12 resolution levels and the four overall robustness metrics. These tables offer the most granular view of model performance, complementing the aggregated analyses in the main text. The abbreviations for sub-tasks are defined as follows: IS (Image Style), ISU (Image Scene Understanding), IA (Instance Attribute), IL (Instance Location), IC (Instance Counting), CIC (Cross-instance Comparison), SRR (Spatial Relation Reasoning), CDR (Chart & Diagram Reasoning), ST (Science & Technology), FC (Function Calculation), Geo (Geometry), SA (Statistical Analysis), DVQA (Doc-oriented VQA), KIE (Key Information Extraction), STU (Scene Text Understanding).
@@ A.3 Details of Res-Bench
To gather a diverse and high-quality candidate pool, we began by aggregating samples from 13 established benchmarks. This included general-purpose MLLM evaluation suites (Chen et al. 2024a; Li et al. 2023), as well as specialized benchmarks targeting specific skills such as visual perception (Fu et al. 2024; Wang et al. 2024b), mathematical reasoning (Lu et al. 2024; Zhang et al. 2024), chart understanding (Masry et al. 2025), spatial reasoning (Wang et al. 2024a), Expert-level multimodal understanding (Yue et al. 2024, 2025), robustness evaluation (Li et al. 2024a) ,and OCR (Liu et al. 2024; Yang et al. 2024).

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) increasingly support dynamic image resolutions. However, current evaluation paradigms primarily assess semantic performance, overlooking the critical question of resolution robustness - whether performance remains stable across varying input resolutions. To address this gap, we introduce Res-Bench, a comprehensive benchmark comprising 14,400 samples across 12 resolution levels and six core capability dimensions. We designed a novel evaluation framework that goes beyond traditional accuracy metrics to capture performance stability. This framework introduces multiple robustness metrics: Spearman’s correlation for assessing resolution-performance trends, and Absolute/Relative Continuous Error (ACE/RCE) for measuring performance volatility. Using these metrics, we conducted a large-scale evaluation of leading MLLMs. Our analysis encompasses: (1) model-centric and task-centric robustness examination, (2) investigation of preprocessing strategies including padding and super-resolution, and (3) exploration of fine-tuning for stability enhancement. Our code is available at https://github.com/cxli2333/Res-Bench.
@@ 1 Introduction
Visual perception and reasoning constitute fundamental components of human perceptual and cognitive capabilities. Notably, humans demonstrate remarkable robustness in processing and interpreting visual information across varying image resolutions. This is largely attributable to the adaptive capabilities of our visual system and the brain’s cognitive completion mechanisms (vision_book; mental_imagery). Specifically, we can adaptively focus on important regions of an image, and for low-resolution images, our brains can fill in missing details based on past experiences and prior knowledge (walsh1998perceptual). The quest to equip machines with such sophisticated perceptual abilities is a central theme in modern AI research. However, while multimodal large language models have demonstrated exceptional capabilities across various domains, it remains unclear how well MLLMs can achieve resolution robustness comparable to human visual systems.
Contemporary advanced MLLMs have incorporated sophisticated designed architectures to enable support for dynamic resolution input. These strategies generally fall into two categories: (1) Native dynamic processing method, such as the Vision Transformer (ViT) (vit) in Qwen2.5-VL (qwen2vl) which employs Multimodal Rotary Position Embedding (MRoPE) and window attention to process images at their native resolution; and (2) Patch-based method, utilized by models like InternVL2.5 (InternVL2.5), LLaVA-OneVision (llava-ov) and LLaVA-UHD (guo2024llava-uhd), which segment high-resolution images into smaller sub-images and thumbnails for individual processing. Current research paradigms predominantly focus on benchmarking model performance at the semantic level, including visual perception (blink), visual question answering (VQA) (vqa), visual reasoning (mathvista), etc.. However, insufficient attention has been directed to the robustness discussion at the visual processing level. For each MLLM, three critical ingredients contribute to the resolution robustness: (1) the processing mechanism for dynamic resolution input entailed in the vision encoder (swin_transformer), (2) the level of question spanning both perceptual and cognitive dimensions (perception_reasoning), and (3) the distribution of training data (survey_mllm_data).
@@
We conduct extensive experiments on Res-Bench to evaluate their robustness to visual inputs. Additionally, for the obtained low-resolution images, we employ data preprocessing methods like white-padding and super-resolution, aiming to validate their effectiveness in both enhancing absolute performance and improving resolution robustness(DBLP:conf/aaai/ZhuWLHL024; DBLP:conf/nips/ZhuZ00HZ24). The evaluation covers multiple proprietary models (GPT-4o( (gpt4o), and Gemini-1.5 Pro (gemini)) and open-source models (e.g.LLaVA-OneVision (llava-ov), Qwen2.5-VL (qwen2vl), mPLUG-Owl3 (mplug), InternVL2.5(InternVL2.5), etc.). We find there is a pervasive lack of resolution robustness. To understand the underlying factors contributing to this instability, we conducted an in-depth analysis of the experimental results, revealing several novel findings that pose new challenges for the development of MLLMs:
@@ 2 Related Work
Multimodal Large Language Model. A fundamental challenge in Multimodal Large Language Models (MLLMs) (qwen2vl; llava-ov; gemini) lies in effectively processing visual information, particularly at high resolutions where fine-grained details are critical. Early models addressed this by resizing images to a low, fixed resolution, a method prone to distorting critical details necessary for tasks like OCR or fine-grained perception (clip; qwenvl). To address these issues, several common solutions have emerged: (1) Native resolution processing methods, such as those used in Qwen2.5-VL (qwen2vl) and Kimi-VL (kimi). (2) Patch-based method: employed by models like LLaVA-OneVision (llava-ov), InternVL2.5 (InternVL2.5), and DeepSeek-VL2 (deepseekvl). While both approaches handle high-resolution inputs, the trade-offs between them regarding resolution robustness remain largely unexplored. This paper systematically investigates the impact of these distinct architectural choices on model stability(DBLP:journals/corr/abs-2507-03657; DBLP:conf/mm/ZhuZTWHZ24).
MLLM Benchmarks. The evaluation of MLLMs has rapidly evolved, moving from single-task benchmarks like VQA (vqa; vizwiz), caption (coco; flickr30k) to comprehensive, multi-faceted evaluation suites. Recent efforts have produced a rich ecosystem of benchmarks targeting a wide array of high-level capabilities, including: Visual perception (blink), Chart reasoning (chartqa; chartqapro; charxiv), Mathematical reasoning (mathvista; mathverse), Spatial reasoning (spatialeval; mvot; bsa), Expert-level multimodal understanding (mmmu; mmmupro), Optical Character Recognition (OCR) capabilities (OCRbench; ccocr). Additionally, comprehensive benchmarks (seed; mmstar) have been developed to evaluate overall multimodal intelligence. While their source images may have varying native resolutions, these benchmarks ignore the impact of the resolution distribution.
Robustness Evaluation. Evaluating the robustness of MLLMs—their ability to maintain stable performance against input perturbations—is a critical area of research. Existing work has primarily focused on several key dimensions: (1) robustness against semantic-level perturbations (MAD-Bench; MMVU); (2) robustness against stylistic or natural variations in images (benchlmm; NaturalBench); and (3) robustness against common image corruptions (robustbench; rbench; zhu2025enhancing). A direct comparison of these approaches, summarized in Table 1, reveals a shared limitation: while they comprehensively examine model stability, they primarily test against content-level corruptions and adversarial inputs. The impact of fundamental visual properties like resolution on model performance remains a critical, under-explored area. We introduce Res-Bench to systematically fill this gap, providing the first dedicated benchmark for evaluating MLLM resolution robustness.
@@ 3.1 Data Collection Process
Preliminary Data Filtering. To establish a comprehensive and high-quality benchmark, we selected 13 existing benchmarks as data sources. These include benchmarks designed for general evaluation of MLLMs (e.g., SEED (seed)) as well as specialized benchmarks for assessing specific capabilities of MLLMs (including AI2D (ai2d), HR-Bench (hrbench), NaturalBench (NaturalBench), SpatialEval (spatialeval), MMMU (mmmu), MMMU-Pro (mmmupro), ChartQA-Pro (chartqapro), BLINK (blink), MathVerse (mathverse), MathVista (mathvista), OCRBench (OCRbench), CCOCR (ccocr)). To ensure the data consists exclusively of high-quality images, we first conducted preliminary filtering by selecting all high-resolution images (with both width and height exceeding 1344 px) as the foundational dataset. Building upon this foundation, we further refined the dataset by leveraging powerful models to select samples that meet visual dependency criteria. Specifically, we employed GPT-4o (gpt4o) and InternVL-2.5-72B (InternVL2.5) as evaluators for this filtering process. We conducted text-only input evaluations by providing textual questions without any accompanying images, while employing circular evaluation to mitigate random biases. A sample was automatically discarded if both models could answer it correctly without the image, as this indicated weak visual dependency. After this filtering process, the number of candidate samples was reduced from 69840 to 2868.
Manual Verification. To further ensure data quality, we performed human expert review on the pre-filtered samples. The evaluation criteria remained consistent with the aforementioned standards. Additionally, we further optimized the data distribution to ensure better balance. To comprehensively evaluate the capabilities of large models across diverse question types and facilitate systematic assessment, we further refined the phrasing of questions for a subset of the data. For instance, we restructured the single-image multi-question format in NaturalBench (NaturalBench) to a single-question multi-image paradigm, while also refining certain question formulations in CCOCR (ccocr). We then systematically generated the full Res-Bench dataset by creating multiple resolution versions for each image, resulting in a total of 14,400 data instances for evaluation. Our data filtering pipeline is illustrated in Figure 2.
@@ 4 Experiment
Baseline. We use Res-Bench to evaluate several leading MLLMs. This includes two powerful proprietary models, GPT-4o (gpt4o) and Gemini 1.5-Pro (gemini), and six powerful open-source models employing different dynamic resolution strategies. Qwen2.5-VL (qwen2vl) , Kimi-VL-Instruct (kimi) directly process native high-resolution images, while InternVL-2.5 (InternVL2.5), LLaVA-OneVision (llava-ov) , MiniCPM-o-2.6 (minicpm) use patch-based method. We also tested model that do not support dynamic resolution for single images, mPLUG-Owl3 (mplug).
@@ 4.2 Analysis of Preprocessing Strategies
Restoring Information Content via Super-Resolution. We investigated whether algorithmically restoring visual details could recover lost performance more effectively. We employed DiffIR (diffir), a classical diffusion-based SR model, to upscale low-resolution images (224px and 448px) to higher resolutions (448px and 896px). Unlike padding, SR aims to intelligently “fill in” missing information, enhancing image clarity. As shown in Table 3, SR-enhanced images significantly outperform their original low-resolution counterparts and also surpass the performance of the simple padded images.
@@ A.2 Experiment Details
This section provides a detailed, per-model breakdown of results on the Res-Bench dataset. We evaluated a diverse suite of eight leading MLLMs, including proprietary models like Gemini 1.5 Pro (gemini) and GPT-4o (gpt4o), as well as open-source models such as Qwen2.5-VL (qwen2vl), LLaVA-OneVision (llava-ov), InternVL2.5 (InternVL2.5), DeepSeek-VL(deepseekvl), MiniCPM-o-2.6 (minicpm) and mPLUG-Owl3 (mplug). For each of the evaluated models, we present the full results across all 15 sub-tasks, including accuracy at each of the 12 resolution levels and the four overall robustness metrics. These tables offer the most granular view of model performance, complementing the aggregated analyses in the main text. The abbreviations for sub-tasks are defined as follows: IS (Image Style), ISU (Image Scene Understanding), IA (Instance Attribute), IL (Instance Location), IC (Instance Counting), CIC (Cross-instance Comparison), SRR (Spatial Relation Reasoning), CDR (Chart & Diagram Reasoning), ST (Science & Technology), FC (Function Calculation), Geo (Geometry), SA (Statistical Analysis), DVQA (Doc-oriented VQA), KIE (Key Information Extraction), STU (Scene Text Understanding).
@@ A.3 Details of Res-Bench
To gather a diverse and high-quality candidate pool, we began by aggregating samples from 13 established benchmarks. This included general-purpose MLLM evaluation suites (mmstar; seed), as well as specialized benchmarks targeting specific skills such as visual perception (blink; hrbench), mathematical reasoning (mathvista; mathverse), chart understanding (chartqapro), spatial reasoning (spatialeval), Expert-level multimodal understanding (mmmu; mmmupro), robustness evaluation (NaturalBench) ,and OCR (OCRbench; ccocr).